OpenStudy (anonymous):

1. How much would 18 ounces of peanut butter cost at $2.24 per pond? 2. A large truck had a load of tons of oranges. How many 5-lb. bags will the oranges make? 3. Mrs. White weighed 83.7 kg. before she began her diet. For the past six months shee lost an average of 2.45 kg. each month. How many did she weigh now? Find her approximate weight in pounds by multipying her weight by 2.2

OpenStudy (anonymous):

16 oz = 1 lb 18 oz * 1/16 * 2.24 = 2.52

OpenStudy (anonymous):

1 ton = 2000 lbs 5*400 = 2000; so 400 bags of orangess

OpenStudy (anonymous):

83.7 - 6(2.45) = 78.8 kg 78.8 * 2.2 = 173.36 lbs

OpenStudy (mertsj):